# Universal Mott quantum criticality in a modified periodic Anderson model

### v1

@@ Abstract
Mott quantum criticality is a central theme in correlated electron physics, observed in systems featuring both continuous zero-temperature transitions and those with finite-temperature critical endpoints. Within dynamical mean-field theory (DMFT), the paradigmatic single-band Hubbard model (SBHM) displays such criticality only above a finite-temperature endpoint. In contrast, the modified periodic Anderson model (MPAM) is a rare example known to host a surface of genuinely continuous Mott quantum critical points (QCPs) at zero temperature. Using DMFT with the numerical renormalization group as an impurity solver, we investigate the finite-temperature, real-frequency properties of the MPAM. Our central finding is the emergence of quantum critical scaling in the electrical resistivity, with critical exponents $z_{\text{met}}=0.76$ and $z_{\text{ins}}=0.66$ on the metallic and insulating sides, respectively. These values fall within the range reported for the SBHM, suggesting that both transitions are governed by a common universality class. We further substantiate the presence of local quantum criticality by demonstrating robust $\omega/T$ scaling in single- and two-particle correlation functions. Finally, we identify novel transport signatures in the optical conductivity, where the distinct evolution of two isosbestic points serves as a unique fingerprint of the QCP. These results establish the MPAM as a canonical model for investigating genuine Mott quantum criticality and support the existence of a universal framework for this fundamental phenomenon.
@@ I Introduction
Quantum phase transitions (QPTs), which occur at absolute zero temperature when a non-thermal control parameter tunes the system between competing ground states, are of fundamental significance in the field of strongly correlated electron systems[1, 2, 3]. A canonical example of such a phenomenon is the interaction-driven metal-to-insulator transition, widely known as the Mott transition[4]. Extensive investigations over recent decades have revealed that this transition can be either first-order [5, 6, 7, 8] or continuous[9, 10, 11]. The single-band Hubbard model (SBHM)[12, 13, 14], the paradigmatic theoretical framework for this problem, correspondingly displays a first-order Mott transition terminating at a finite-temperature critical point[15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Intriguingly, theoretical analyses of the SBHM have shown a characteristic quantum critical scaling in the electrical resistivity above the critical temperature[23, 25, 24], a feature generally associated with quantum criticality. This scaling behavior was subsequently realized in experiments on materials with first-order transitions[8], yielding a comparable scaling exponent ($\nu z$). Remarkably, a similar collapse with a consistent scaling exponent has also been observed in Moiré superlattices[10], which exhibit a continuous Mott transition. This convergence implies that the critical exponents derived from the SBHM may be universal, applicable even to systems that undergo a continuous Mott transition. While continuous Mott transitions have been theoretically realized in lower-dimensional or geometrically frustrated Hubbard models, such as on the 1/5-depleted square lattice[26], a comprehensive analysis of their critical exponents is often not performed. The only other model known to exhibit a continuous ($T_{c}$=0) Mott transition in higher dimensions is a modified periodic Anderson model (MPAM)[27, 28], which describes a continuous Mott transition within the paramagnetic phase.
@@
The MPAM [27, 28] exhibits a quantum phase transition within the paramagnetic regime, separating a Fermi-liquid metal from a Mott insulator. At the quantum critical point (QCP), a singular density of states emerges, characterized by the formation of a soft gap[28]. This soft gap form of the hybridization is incorporated in the pseudogap Anderson model[29, 30, 31, 32, 33], which results in an impurity quantum phase transition, which is otherwise absent. In contrast, within the MPAM, this form emerges naturally within DMFT at the QCP[28].
The soft-gap spectral function of the MPAM at the QCP corresponds to the local moment fixed point of the pseudogap Anderson model. This form of the spectral function has been reported in the SBHM [24] at the lower edge of the co-existent region ($U=U_{c1}$).
@@
While previous studies[27, 28] have provided significant insights, the absence of finite-temperature real-frequency data leaves several fundamental questions unresolved. These include: Does the MPAM display Fermi-liquid scaling in its spectral function? How does the QCP influence the adjacent phases, as reflected in the correlation functions? Does the MPAM exhibit $\omega/T$ scaling, a hallmark of local quantum criticality? What are the key transport signatures of a QCP? Does the resistivity exhibit a scaling collapse analogous to that observed in the SBHM, thereby suggesting a common universality class for Mott quantum criticality? To address these questions, we employ the numerical renormalization group (NRG) method within the dynamical mean-field theory (DMFT) framework [19] to study the MPAM directly at finite temperatures and on the real-frequency axis.
@@ II MODEL AND FORMALISM
The MPAM in the particle–hole symmetric limit has been examined within DMFT in the paramagnetic regime, using the continuous-time quantum Monte Carlo (CT-QMC) method at finite temperatures [28] and the local moment approach (LMA) at zero temperature [27] as impurity solvers. Both studies demonstrate that the model exhibits a surface of QCPs in the $U-t_{\perp}$ plane, separating the Fermi liquid (FL) phase from the Mott insulating phase. Throughout this work, we fix the hybridization at $V=0.44$ and restrict ourselves to the paramagnetic, particle-hole symmetric phase, which corresponds to setting $\epsilon_{c}=\epsilon_{M}=0$ and $\epsilon_{f}=-U/2$. The chemical potential is set to zero throughout.
In this work, we solve the lattice model within the framework of dynamical mean-field theory (DMFT). DMFT is a powerful non-perturbative approach that maps a lattice problem onto an effective single-impurity model, which is then solved self-consistently [34, 19, 35]. This mapping becomes exact in the limit of infinite spatial dimensions and neglects all nonlocal contributions to the retarded self-energy. Consequently, the retarded self-energy is purely local, i.e., $\Sigma_{\text{f}}(\omega,\mathbf{k})\equiv\Sigma_{\text{f}}(\omega)$. The self-consistency condition for the MPAM within the DMFT framework is expressed as
@@
Here, retarded $f$-Green’s function is given as,
@@
where,
@@
$G_{\text{cc}}(\omega,\epsilon)\equiv\langle\langle c_{\textbf{k}\sigma};c_{\textbf{k}\sigma}\rangle\rangle_{\omega},\,G_{\text{mm}}(\omega,\epsilon)\equiv\langle\langle c_{M\textbf{k}\sigma};c_{M\textbf{k}\sigma}\rangle\rangle_{\omega},\,$
@@
$G_{\text{mc}}(\omega,\epsilon)\equiv\langle\langle c_{M\textbf{k}\sigma};c_{\textbf{k}\sigma}\rangle\rangle_{\omega},G_{\text{cm}}(\omega,\epsilon)\equiv\langle\langle c_{\textbf{k}\sigma};c_{M\textbf{k}\sigma}\rangle\rangle_{\omega}.$
@@
To satisfy the equality in Eq. 2, the impurity problem must be solved iteratively using the self-consistently determined hybridization function at each step. The impurity problem can be solved using a variety of methods, including Iterative Perturbation Theory (IPT) [36, 37, 38], the numerically exact Continuous-Time Quantum Monte Carlo (CTQMC) method [39, 40, 41, 42], the Local Moment Approach (LMA) [43, 44, 45], and the Numerical Renormalization Group (NRG) [46]. Each of the methods has its own drawbacks. For example, IPT is a second order approximation theory and fails to capture non-perturbative effects. CTQMC is numerically exact but gives results on the Matsubara axis. To obtain real frequency data, it is necessary to perform analytic continuation, an ill-defined procedure. Furthermore, CTQMC calculations get prohibitively expensive, particularly at very low temperatures. On the other hand, NRG is a very powerful tool and a non-perturbative technique that works in real frequencies. To address the impurity problem within DMFT, we employ the NRG method, utilizing the open-source implementation developed by Rok Žitko [47].
@@
The NRG, a non-perturbative technique, was originally developed by Wilson [46]. To accurately capture all energy scales, the conduction bath in NRG is discretized into logarithmic intervals $\Lambda^{n+1},\Lambda^{n}$}] ($n=0,1,2,\ldots$), where $\Lambda>1$ is the discretization parameter. The width of each interval is given by $d_{n}=\Lambda^{-n}(1-\Lambda^{-1})$, and the continuum limit is recovered as $\Lambda\rightarrow 1$. The continuum of conduction band states is then replaced by a discrete set of states, which is mapped onto a semi-infinite tight-binding chain with the impurity located at one end. The hopping amplitudes along the chain decay as $\Lambda^{-n/2}$. The resulting Hamiltonian is solved iteratively, but the exponential growth of the Hilbert space with system size presents a computational challenge. To address this, Wilson introduced a truncation scheme[46] in which the number of retained states is limited to $N_{\text{states}}$, as the discarded high-energy states have negligible influence on the low-energy physics. NRG has proven highly successful in the study of quantum impurity problems [48, 49, 50, 51, 52], and has also been widely employed as an impurity solver within DMFT for lattice models such as the Hubbard model [53, 20], the PAM [54], and others. In this work, we adopt a discretization parameter of $\Lambda=2$ and perform twist averaging over $N_{z}=16$ [55], whereby separate calculations are carried out for different interleaved discretization grids, and the resulting quantities are subsequently averaged. To accelerate the convergence of DMFT iterations, we employ Broyden’s mixing scheme [56].
@@ III.1 Analysis of single particle quantities
In this section, we demonstrate universal FL scaling, the continuous vanishing of the quasiparticle weight and Mott gap at the QCP, and identify the crossover scale to the quantum critical region from the self-energy. Understanding universal behavior across different phases is crucial, as it provides deeper insight into the nature of the phases and facilitates comparison with experimental results [57]. In the FL regime of the MPAM, and in the limit $\omega\rightarrow 0$, the single-particle $f$-spectral function scales as $\text{A}_{\text{f}}(\omega)=-\text{Im}[G_{\text{loc}}(\omega)]/\pi=(Z^{2}V^{2}/\omega^{2})\text{A}_{\text{c}}(\omega)$ [27], where $Z=\left(1-\frac{d}{d\omega}\text{Re}\,\Sigma_{\text{f}}(\omega)\right)^{-1}$ denotes the quasiparticle weight and $\text{A}_{\text{c}}(\omega)=\text{-Im}[G_{\text{c}}(\omega)]$ $\left(G_{\text{c}}(\omega)=\int\,d\epsilon\,G_{\text{c}}(\omega,\epsilon)\right)$ is the $c$-band spectral function. The $c$-band spectral function can be given as,
@@
Figure 1: (Top) Rescaled spectral weight, $\text{A}_{\text{f}}(\omega)V^{2}/t_{\perp}^{2}$, as a function of $\omega t_{\perp}^{2}/ZV^{2}$ at $T=10^{-3}$. (Bottom) Energy scales: quasi-particle weight (Z{Z})(solid lines- red), crossover scale(dashed-blue) and Mott gap ($\Delta_{g}$)(solid lines- black) vs. inter-orbital coupling $t_{\perp}$. The crossover scale is extracted from the self-energy deviation from the power-law form(see text corresponding to Fig. 2). $Z$ and $\Delta_{g}$, calculated at $T=0$, vanish with exponents $2.0$ and $1.05$. Inset: $\text{A}_{\text{f}}(\omega)$ for different $t_{\perp}$. Parameters used: $U=1.75$.
@@
The top panel of Fig. 1 displays the scaled spectral function in the FL regime. As expected, the spectral function exhibits a scaling collapse as $\omega\rightarrow 0$ and remains pinned at $\omega=0$. This observation confirms the principle of adiabatic continuity, a necessary condition for the existence of a Fermi liquid phase. In the Mott insulating phase, the Hubbard bands move further apart with increasing $t_{\perp}$ (see the inset of the bottom panel in Fig. 1). To quantify the energy scales, we plot both the quasiparticle weight $Z$ and the Mott gap $\Delta_{g}$ calculated in the limit $T\rightarrow 0$ as functions of the inter-orbital coupling $t_{\perp}$ (see the bottom panel of Fig. 1). For $t_{\perp}<t_{\perp,c}$, the quasiparticle weight vanishes continuously as $t_{\perp}\rightarrow t_{\perp,c}^{-}$ with an exponent of approximately $2.0$, which is close to the DMFT+LMA[27] result of $\approx 2.5$. In the insulating phase, the Mott gap $\Delta_{g}$ vanishes at the transition with an exponent of $\approx 1.05$, which is in good agreement with the value $\approx 1.0$ obtained using DMFT+LMA [27].
@@
Our calculated exponents, while consistent with DMFT+LMA [27] results, differ from those obtained using DMFT+CTQMC [28]. This discrepancy can be attributed to the distinct methodologies for determining the quasiparticle weight, $Z$. The previous CTQMC study approximated $Z$ from the first Matsubara frequency point, with this value then extrapolated to $T=0$. In contrast, our approach leverages the NRG solver to compute the real-frequency self-energy. This enables a direct calculation of $Z$ from its derivative at zero frequency. Beyond low-energy scales, we investigate whether an additional scale characterizes the quantum critical region. To explore this, we analyze the self-energy.
@@
The imaginary part of the self-energy in the FL phase varies as $(\omega^{2}+\pi^{2}T^{2})$. In contrast, in the Mott insulating (MI) phase, $\text{Im}\,\Sigma_{\text{f}}(\omega)$ diverges. Due to scale invariance at the QCP, $\text{Im}\,\Sigma_{\text{f}}(\omega)$ also diverges at criticality. Fig. 2 shows the plot of $\text{Im}\,\Sigma_{\text{f}}(\omega)$ for various values of $t_{\perp}$. In the FL regime ($t_{\perp}\ll t_{\perp,c}$), $\text{Im}\,\Sigma_{\text{f}}(\omega)$ exhibits a quadratic behavior in $\omega$, while in the MI regime ($t_{\perp}\gg t_{\perp,c}$), it diverges. At the critical point $t_{\perp,c}=1.033$, $\text{Im}\,\Sigma_{\text{f}}(\omega)$ follows a power-law behavior with a diverging exponent of 0.35-0.35, as found in Ref. 28. As we move away from the QCP, remnants of the power-law behavior are still visible in both the FL and MI regions. We identify the frequency at which $\text{Im}\,\Sigma_{\text{f}}(\omega)$ begins to deviate from the power-law form on both sides as a crossover scale. The region between these crossover scales represents the quantum critical region. This scale is plotted as a function of $t_{\perp}$ in the inset of Fig. 1. The crossover frequency scale vanishes as the system approaches the QCP. The proximity of the crossover energy scale to the Mott gap on the insulating side of the transition indicates the absence of any intermediate phases between the quantum critical region and the Mott insulator. However, in the FL region, the crossover scale is well separated from the Fermi liquid scale, Z{Z}. This indicates that there is a bad metal region between the quantum critical region and the FL region, similar to the one found in the SBHM[59]. In the next section, we examine the quantum critical region and its scaling behavior.

### v2

@@ Abstract
Mott quantum criticality is a central theme in correlated electron physics, with comparable critical exponents observed in materials featuring both continuous zero-temperature transitions and those with finite-temperature critical endpoints. Such criticality was first predicted theoretically for the single-band Hubbard model (SBHM). Within dynamical mean-field theory (DMFT), the SBHM undergoes a first-order transition at $T=0$ but displays quantum critical scaling above its finite-temperature critical point. However, a comprehensive analysis of a system exhibiting a continuous Mott transition at zero temperature has been lacking. To this end, the modified periodic Anderson model (MPAM) is a rare example known to host a surface of continuous Mott quantum critical points (QCPs). While previous studies of the MPAM characterized its QCP and showed the emergence of a pseudogap Anderson model at the QCP, the analysis was restricted to the Matsubara frequency axis, leaving key questions unresolved: What are the signatures of Mott quantum criticality in transport properties, and do the critical exponents align with the universal behavior seen in the SBHM and experiments? To address these questions, we employ DMFT with the numerical renormalization group as an impurity solver to investigate the real-frequency properties of the MPAM. Our central finding is the emergence of quantum critical scaling in the electrical resistivity, with exponents $\nu z_{\text{met}}\approx 0.76$ and $\nu z_{\text{ins}}\approx 0.66$ on the metallic and insulating sides, respectively. These values fall within the range reported for the SBHM and observed in experiments, suggesting that both transitions are governed by a common universality class. We further substantiate the presence of local quantum criticality by demonstrating robust $\omega/T$ scaling in single- and two-particle correlation functions. Finally, we identify novel signatures in the optical conductivity, where the distinct evolution of two isosbestic points provides a unique fingerprint of Mott quantum criticality. These results establish the MPAM as a canonical model for investigating Mott quantum criticality and support the existence of a universal framework for this fundamental phenomenon.
@@ I Introduction
Quantum phase transitions (QPTs), which occur at absolute zero temperature when a non-thermal control parameter tunes the system between competing ground states, are of fundamental significance in the field of strongly correlated electron systems [sondhi1997continuous, sachdev1999quantum, vojta2003quantum]. A canonical example of such a phenomenon is the interaction-driven metal-to-insulator transition, widely known as the Mott transition [imada1998metal]. Extensive investigations over recent decades have revealed that this transition can be either first-order  [mcwhan1973metal, yao1996electrical, limelette2003mott, furukawa2015quantum] or continuous [neumann2007bilayer, li2021continuous, zheng2022quantum]. The single-band Hubbard model (SBHM) [hubbard1963electron, gutzwiller1963effect, kanamori1963electron], the paradigmatic theoretical framework for this problem, correspondingly displays a first-order Mott transition terminating at a finite-temperature critical point [rozenberg1992mott, jarrell1992hubbard, georges1992numerical, rozenberg1994mott, georges1996dynamical, bulla1999zero, bulla2001finite, werner2007doping, terletska2011quantum, eisenlohr2019mott]. Intriguingly, theoretical analyses of the SBHM have shown a characteristic quantum critical scaling in the electrical resistivity above the critical temperature [terletska2011quantum, vuvcivcevic2013finite, eisenlohr2019mott], a feature generally associated with quantum criticality. This scaling behavior was subsequently realized in experiments on materials with first-order transitions [furukawa2015quantum], yielding a comparable scaling exponent ($\nu z$). Remarkably, a similar collapse with a consistent scaling exponent has also been observed in Moiré superlattices [li2021continuous], which exhibit a continuous Mott transition. This convergence implies that the critical exponents derived from the SBHM may be universal, applicable even to systems that undergo a continuous Mott transition. While continuous Mott transitions have been theoretically realized in lower-dimensional or geometrically frustrated Hubbard models, such as on the 1/5-depleted square lattice [yanagi2014continuous], a comprehensive analysis of their critical exponents is often not performed. The only other model known to exhibit a continuous ($T_{c}$=0) Mott transition in higher dimensions is a modified periodic Anderson model (MPAM) [sen2016quantum, kksujan2023emergent], which describes a continuous Mott transition within the paramagnetic phase.
@@
The MPAM is a periodic Anderson model (PAM) [georges1996dynamical, smith2003kondo, vidhyadhiraja2004dynamics, logan2016mott] coupled to an additional conduction band via an inter-orbital coupling $t_{\perp}$. In the particle-hole symmetric limit, the PAM is a Kondo insulator [smith2003kondo, logan2016mott]. Adding a conduction band to a PAM induces a finite spectral weight in the erstwhile localized f-band for any non-zero $t_{\perp}$ [sen2016quantum]. Consequently, a finite interaction strength is required to drive a metal-to-insulator transition [sen2016quantum]. Indeed, the interacting MPAM [sen2016quantum, kksujan2023emergent] exhibits a quantum phase transition within the paramagnetic regime, separating a Fermi-liquid metal from a Mott insulator. At the quantum critical point (QCP), a singular density of states emerges, characterized by the formation of a soft gap [kksujan2023emergent]. This soft gap form of the hybridization is incorporated in the pseudogap Anderson model [withoff1990phase, fritz2004phase, glossop2005local, glossop2011critical, chowdhury2015critical], which results in an impurity quantum phase transition, which is otherwise absent. In contrast, within the MPAM, this form emerges naturally within DMFT at the QCP [kksujan2023emergent]. The soft-gap spectral function of the MPAM at the QCP corresponds to the local moment fixed point of the pseudogap Anderson model. This form of the spectral function has been reported in the SBHM [eisenlohr2019mott] at the lower edge of the co-existence region ($U=U_{c1}$).
@@
While previous studies [peters2013kondo, sen2015spectral, sen2016quantum, hu2017effects, kksujan2023emergent, majumder2025mott] have provided significant insights, the absence of finite-temperature real-frequency data leaves several fundamental questions unresolved. These include: Does the MPAM display Fermi-liquid scaling in its spectral function? How does the QCP influence the adjacent phases, as reflected in the correlation functions? Does the MPAM exhibit $\omega/T$ scaling, a hallmark of local quantum criticality? What are the key transport signatures of Mott quantum criticality? Does the resistivity exhibit a scaling collapse analogous to that observed in the SBHM, thereby suggesting a common universality class for Mott quantum criticality? To address these questions, we employ the numerical renormalization group (NRG) method within the dynamical mean-field theory (DMFT) framework [georges1996dynamical] to study the MPAM directly at finite temperatures and on the real-frequency axis.
@@ II MODEL AND FORMALISM
The MPAM in the particle–hole symmetric limit has been examined within DMFT in the paramagnetic regime, using the continuous-time quantum Monte Carlo (CT-QMC) method at finite temperatures  [kksujan2023emergent] and the local moment approach (LMA) at zero temperature  [sen2016quantum] as impurity solvers. Both studies demonstrate that the model exhibits a surface of QCPs in the $U-t_{\perp}$ plane, separating the Fermi liquid (FL) phase from the Mott insulating phase. Throughout this work, we fix the hybridization at $V=0.44$ and restrict ourselves to the paramagnetic, particle-hole symmetric phase, which corresponds to setting $\epsilon_{c}=\epsilon_{M}=0$ and $\epsilon_{f}=-U/2$. The chemical potential is set to zero throughout.
In this work, we solve the lattice model within the framework of dynamical mean-field theory (DMFT). DMFT is a powerful non-perturbative approach that maps a lattice problem onto an effective single-impurity model, which is then solved self-consistently  [metzner1989correlated, georges1996dynamical, vollhardt2011dynamical]. This mapping becomes exact in the limit of infinite spatial dimensions and neglects all nonlocal contributions to the retarded self-energy. Consequently, the retarded self-energy is purely local, i.e., $\Sigma_{\text{f}}(\omega,\mathbf{k})\equiv\Sigma_{\text{f}}(\omega)$. The self-consistency condition for the MPAM within the DMFT framework is expressed as
@@
Here, the retarded $f$-Green’s function is given as,
@@
Here, the Green’s functions are defined in standard Zubarev [zubarev1960double, vidhyadhiraja2007specific] notation as:
@@
$G_{\text{cc}}(\omega,\epsilon)\equiv\langle\langle c_{\textbf{k}\sigma};c_{\textbf{k}\sigma}^{\dagger}\rangle\rangle_{\omega},\,G_{\text{mm}}(\omega,\epsilon)\equiv\langle\langle c_{M\textbf{k}\sigma};c_{M\textbf{k}\sigma}^{\dagger}\rangle\rangle_{\omega},$
@@
$G_{\text{mc}}(\omega,\epsilon)\equiv\langle\langle c_{M\textbf{k}\sigma};c_{\textbf{k}\sigma}^{\dagger}\rangle\rangle_{\omega},G_{\text{cm}}(\omega,\epsilon)\equiv\langle\langle c_{\textbf{k}\sigma};c_{M\textbf{k}\sigma}^{\dagger}\rangle\rangle_{\omega}.$
@@
To satisfy the equality in Eq. 2, the impurity problem must be solved iteratively using the self-consistently determined hybridization function at each step. The impurity problem can be solved using a variety of methods, including Iterative Perturbation Theory (IPT) [yamada1975, yosida1975perturbation, kajueter1996new], the numerically exact Continuous-Time Quantum Monte Carlo (CTQMC) method [rubtsov2004continuous, rubtsov2005continuous, werner2006hybridization, gull2008continuous], the Local Moment Approach (LMA) [logan1998local, vidhyadhiraja2004dynamics, galpin2005single], and the Numerical Renormalization Group (NRG) [wilson1975renormalization]. Each of the methods has its own drawbacks. For example, IPT is a second order approximation theory and fails to capture non-perturbative effects. CTQMC is numerically exact but gives results on the Matsubara axis. To obtain real frequency data, it is necessary to perform analytic continuation, an ill-defined procedure. Furthermore, CTQMC calculations get prohibitively expensive, particularly at very low temperatures. On the other hand, NRG is a non-perturbative technique that works in real frequencies. To address the impurity problem within DMFT, we employ the NRG method, utilizing the open-source implementation developed by Rok Žitko  [vzitko2009energy].
@@
The NRG, a non-perturbative technique, was originally developed by Wilson  [wilson1975renormalization]. To accurately capture all energy scales, the conduction bath in NRG is discretized into logarithmic intervals $\Lambda^{n+1},\Lambda^{n}$}] ($n=0,1,2,\ldots$), where $\Lambda>1$ is the discretization parameter. The width of each interval is given by $d_{n}=\Lambda^{-n}(1-\Lambda^{-1})$, and the continuum limit is recovered as $\Lambda\rightarrow 1$. The continuum of conduction band states is then replaced by a discrete set of states, which is mapped onto a semi-infinite tight-binding chain with the impurity located at one end. The hopping amplitudes along the chain decay as $\Lambda^{-n/2}$. The resulting Hamiltonian is solved iteratively, but the exponential growth of the Hilbert space with system size presents a computational challenge. To address this, Wilson introduced a truncation scheme [wilson1975renormalization] in which the number of retained states is limited to $N_{\text{states}}$, as the discarded high-energy states have negligible influence on the low-energy physics. The NRG has proven highly successful in the study of quantum impurity problems  [krishna1980renormalization, sakai1989single, costi1994transport, bulla1998numerical, bulla2008numerical], and has also been widely employed as an impurity solver within DMFT for lattice models such as the Hubbard model  [sakai1994application, bulla1999zero], the PAM  [pruschke2000low], and others. In this work, we adopt a discretization parameter of $\Lambda=2$ and perform twist averaging over $N_{z}=16$  [campo2005alternative], whereby separate calculations are carried out for different interleaved discretization grids, and the resulting quantities are subsequently averaged. To accelerate the convergence of DMFT iterations, we employ Broyden’s mixing scheme  [vzitko2009convergence].
@@ III.1 Analysis of single particle quantities
In this section, we demonstrate universal FL scaling, the continuous vanishing of the quasiparticle weight and Mott gap at the QCP, and identify the crossover scale to the quantum critical region from the self-energy. Understanding universal behavior across different phases is crucial, as it provides deeper insight into the nature of the phases and facilitates comparison with experimental results [vidhyadhiraja2005optical]. In the FL regime of the MPAM, and in the limit $\omega\rightarrow 0$, the single-particle $f$-spectral function scales as $\text{A}_{\text{f}}(\omega)=-\text{Im}[G_{\text{loc}}(\omega)]/\pi=(Z^{2}V^{2}/\omega^{2})\text{A}_{\text{c}}(\omega)$ [sen2016quantum], where $Z=\left(1-\frac{d}{d\omega}\text{Re}\,\Sigma_{\text{f}}(\omega)\right)^{-1}$ denotes the quasiparticle weight and $\text{A}_{\text{c}}(\omega)=\text{-Im}[G_{\text{c}}(\omega)]$ $\left(G_{\text{c}}(\omega)=\int\,d\epsilon\,G_{\text{c}}(\omega,\epsilon)\right)$ is the $c$-band spectral function. In the low-frequency limit ($\omega\rightarrow 0$), the c-band spectral function (see Appendix B for the derivation) is given by:
@@
Figure 1: (Top) Rescaled spectral weight, $\text{A}_{\text{f}}(\omega)V^{2}/t_{\perp}^{2}$, as a function of $\omega t_{\perp}^{2}/ZV^{2}$ at $T=10^{-3}$. (Bottom) Energy scales: quasi-particle weight (Z{Z})(solid lines- red), crossover scale(dashed-blue) and Mott gap ($\Delta_{g}$)(solid lines- black) vs. inter-orbital coupling $t_{\perp}$. The crossover scale is extracted from the self-energy deviation from the power-law form(see text corresponding to Fig.  2). $Z$ and $\Delta_{g}$, calculated at $T=0$, vanish with exponents $2.0$ and $1.05$. Inset: $\text{A}_{\text{f}}(\omega)$ for different $t_{\perp}$. Parameters used: $U=1.75$.
@@
The top panel of Fig. 1 displays the scaled spectral function in the FL regime. As expected, the spectral function exhibits a scaling collapse as $\omega\rightarrow 0$ and remains pinned at $\omega=0$. This observation confirms the principle of adiabatic continuity, a necessary condition for the existence of a Fermi liquid phase. In the Mott insulating phase, the Hubbard bands move further apart with increasing $t_{\perp}$ (see the inset of the bottom panel in Fig. 1). To quantify the energy scales, we plot both the quasiparticle weight $Z$ and the Mott gap $\Delta_{g}$ calculated in the limit $T\rightarrow 0$ as functions of the inter-orbital coupling $t_{\perp}$ (see the bottom panel of Fig. 1). For $t_{\perp}<t_{\perp,c}$, the quasiparticle weight vanishes continuously as $t_{\perp}\rightarrow t_{\perp,c}^{-}$ with an exponent of approximately $2.0$, which is close to the DMFT+LMA [sen2016quantum] result of $\approx 2.5$. In the insulating phase, the Mott gap $\Delta_{g}$ vanishes at the transition with an exponent of $\approx 1.05$, which is in good agreement with the value $\approx 1.0$ obtained using DMFT+LMA [sen2016quantum].
@@
Our calculated exponents, while consistent with DMFT+LMA [sen2016quantum] results, differ from those obtained using DMFT+CTQMC [kksujan2023emergent]. This discrepancy can be attributed to the distinct methodologies for determining the quasiparticle weight, $Z$. The previous CTQMC study approximated $Z$ from the first Matsubara frequency point, with this value then extrapolated to $T=0$. In contrast, our approach leverages the NRG solver to compute the real-frequency self-energy. This enables a direct calculation of $Z$ from its derivative at zero frequency. Beyond low-energy scales, we investigate whether an additional scale characterizes the quantum critical region. To explore this, we analyze the self-energy.
@@
The imaginary part of the self-energy in the FL phase varies as $(\omega^{2}+\pi^{2}T^{2})$. In contrast, in the Mott insulating (MI) phase, $\text{Im}\,\Sigma_{\text{f}}(\omega)$ diverges. Due to scale invariance at the QCP, $\text{Im}\,\Sigma_{\text{f}}(\omega)$ also diverges at criticality. Fig. 2 shows the plot of $\text{Im}\,\Sigma_{\text{f}}(\omega)$ for various values of $t_{\perp}$. In the FL regime ($t_{\perp}\ll t_{\perp,c}$), $\text{Im}\,\Sigma_{\text{f}}(\omega)$ exhibits a quadratic behavior in $\omega$, while in the MI regime ($t_{\perp}\gg t_{\perp,c}$), it diverges. At the critical point $t_{\perp,c}=1.033$, $\text{Im}\,\Sigma_{\text{f}}(\omega)$ follows a power-law behavior with a diverging exponent of 0.35-0.35, as found in Ref. kksujan2023emergent. As we move away from the QCP, remnants of the power-law behavior remain visible in both the FL and MI regions. We identify the frequency at which $\text{Im}\,\Sigma_{\text{f}}(\omega)$ begins to deviate from the power-law form on both sides as a crossover scale. The region between these crossover scales represents the quantum critical region. This scale is plotted as a function of $t_{\perp}$ in the bottom panel of Fig. 1. The crossover frequency scale vanishes as the system approaches the QCP. The proximity of the crossover energy scale to the Mott gap on the insulating side of the transition indicates the absence of any intermediate phases between the quantum critical region and the Mott insulator. However, in the FL region, the crossover scale is well separated from the Fermi liquid scale, Z{Z}. This indicates that there is a bad metal region between the quantum critical region and the FL region, similar to the one found in the SBHM [dasari2017quantum]. In the next section, we examine the quantum critical region and its scaling behavior.